OpenStudy (anonymous):

Jim has grades of 84, 65, and 76 on three math tests. What grade must he obtain on the next test to have an average of exactly 80 for the four tests? A. 87 B. 95 C. 85 D. 92

OpenStudy (anonymous):

(84 + 65 + 76 + x) / 4 = 80 can you solve it from here ??

OpenStudy (here_to_help15):

Try checking all the answers. 84+65+76=225 A. 225+87=312 312/4=78 B. 225+95=320 320/4=80 C. 225+85=310 310/4=77.5 D. 225+92=317 317/4=79.25 The answer is B. 95

OpenStudy (anonymous):

(225 + x) / 4 = 80 -- multiply by 4 on both sides 225 + x = 320 x = 320 - 225 x = 95....yep...Here_to_Help is correct